OpenStudy (anonymous):

Simplify the expression 3x^-4y^5 / (2x^3y^-7)^-2

OpenStudy (alana242):

3x^-4y^5 / (2x^3y^-7)^-2 = 3 x^-4 y^5 / (2^-2 x^-6 y^14) = 3(y^5)(2^2)(x^6) / (y^14)(x^4) = 12x^2 / y^9

OpenStudy (anonymous):

thanks!!! how did you make it positive?

OpenStudy (alana242):

Np...Well I Cant Really Explain